given market demand Q=100-P if the market supply function for the smaller firms is given by S=0.5P. And the cost function of the dominant firms is TC=10+40QD, where QD= output of the dominant firm. a) Find the market price and output of the dominant firm at equilibrium. b) Find the output level to be supplied by the smaller firms.

1
Expert's answer
2022-01-12T09:11:10-0500

"a)"

Market price and output of the dominant firm:

At equilibrium: market demand = market supply

Therefore,

100 – P = 0.5P

100 = 0.5P + P

100 = 1.5P

"P=\frac{100}{1.5} = 66.67"


Given P=66.67, we equate for Q

Q = 100 – P

Q = 100 – 66.67

Q = 33.33

Therefore, the market price is 66.67 and the market quantity for the dominant firm is 33.33units


"b)"

Output level to be supplied by the smaller firms:

Smaller firms will supply output at the equilibrium price

The Equilibrium price, "P= 66.67" and the smaller firms supply function is,

"S=0.5P"

Therefore,

 "S=0.5(66.67) = 33.33 units"

Thus, smaller firms will supply an output level of 33.33 units.